OpenStudy (anonymous):

Pleaseee Helpppp What is the simplified form of the seventh root of x times the seventh root of x times the seventh root of x times the seventh root of x?

OpenStudy (anonymous):

\[\sqrt[7]{x} *\sqrt[7]{x}*\sqrt[7]{x} *\sqrt[7]{x} = x ^{1/7}*x ^{1/7}*x ^{1/7}*x ^{1/7}\] \[x ^{1/7}*x ^{1/7}*x ^{1/7}*x ^{1/7} = x^{1/7+1/7+1/7+1/7}\]

OpenStudy (anonymous):

Thankss, could you help me with some more questions?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Frank and Gertrude mow lawns for extra money over the summer. Frank's income is determined by f(x) = 5x + 15, where x is the number of hours. Gertrude's income is g(x) = 4x + 20. If Frank and Gertrude were to combine their efforts, their income would be h(x) = f(x) + g(x). Assume Frank works 4 hours. Create the function h(x), and indicate if Frank will make more money working alone or by teaming with Gertrude. h(x) = x + 5, work alone h(x) = x + 5, team with Gertrude h(x) = 9x + 35, work alone h(x) = 9x + 35, team with Gertrude

OpenStudy (anonymous):

f(x) = 5x + 15, g(x) = 4x + 20 h(x) = f(x) + g(x) h(x) =???

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

in which part?

OpenStudy (anonymous):

5x+15+4x+20 right?

OpenStudy (anonymous):

right. join the same variables

OpenStudy (anonymous):

9x+35 Right?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

now if x = 4 where do you think frank will earn more? in f(x) or in h(x)

OpenStudy (anonymous):

h(x)

OpenStudy (anonymous):

correct. and when will he earn more if he is working with Gertrude, where he has to share the earnings, or if he is working alone.

OpenStudy (anonymous):

With Gertrude and 35?

OpenStudy (anonymous):

@sangya21 ?

OpenStudy (anonymous):

I think alone cause if with Gertrude, he will get 35.1 and alone he will get 71 9*4 +35 = 36+35 = 71

OpenStudy (anonymous):

h(x) = 9x + 35, work alone so that's the answer?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Okayy, I'm going to go with that.

OpenStudy (anonymous):

What is the 1st term when 1 + 6x2 + 5x4 + 3x3 is arranged in descending order? Could you help with this one?

OpenStudy (anonymous):

is this addition sign between them?

OpenStudy (anonymous):

Yeah I think

OpenStudy (anonymous):

5x^4 + 3x^3 + 6x^2 +1 if x is positive and reverse the above eq. id x is -ve

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

when x is +ve, the number increases with increase in power and when x is -ve, the number decreases with increase in power

OpenStudy (anonymous):

What is +ve and -ve?

OpenStudy (anonymous):

Positive and negative

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

These are the answer choices 1 6x2 5x4 3x3

OpenStudy (anonymous):

5x^4 + 3x^3 + 6x^2 +1 lets consider x is positive

OpenStudy (anonymous):

I'm not sure

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

I can suggest someone else for you then .